# Diffusion Dynamics Models with Generative State Estimation for Cloth Manipulation

### v1

@@ Abstract
Manipulating deformable objects like cloth is challenging due to their complex dynamics, near-infinite degrees of freedom, and frequent self-occlusions, which complicate state estimation and dynamics modeling. Prior work has struggled with robust cloth state estimation, while dynamics models, primarily based on Graph Neural Networks (GNNs), are limited by their locality. Inspired by recent advances in generative models, we hypothesize that these expressive models can effectively capture intricate cloth configurations and deformation patterns from data. Building on this insight, we propose a diffusion-based generative approach for both perception and dynamics modeling. Specifically, we formulate state estimation as reconstructing the full cloth state from sparse RGB-D observations conditioned on a canonical cloth mesh and dynamics modeling as predicting future states given the current state and robot actions. Leveraging a transformer-based diffusion model, our method achieves high-fidelity state reconstruction while reducing long-horizon dynamics prediction errors by an order of magnitude compared to GNN-based approaches. Integrated with model-predictive control (MPC), our framework successfully executes cloth folding on a real robotic system, demonstrating the potential of generative models for manipulation tasks with partial observability and complex dynamics.
Figure 1: We present UniClothDiff, a unified diffusion-based framework for cloth manipulation. Our approach integrates state estimation and dynamics modeling under a consistent architecture and training paradigm. Our diffusion-based perception model generates cloth states from partial observations, and the diffusion-based dynamics model generates physically plausible future states conditioned on action sequences, enabling robust model-based control. Our work demonstrates the potential of diffusion models in state estimation and dynamics modeling for manipulation tasks involving partial observability and complex dynamics.
I Introduction
@@
Textile deformable objects, such as clothing, are ubiquitous in daily life, with applications ranging from healthcare and domestic tasks to the textile industry. However, manipulating deformable objects remains a persistent challenge in robotics [20, 42]. Effective handling of cloth requires a robot to estimate its geometric structure despite self-occlusions and its infinite-dimensional state space. Furthermore, optimizing actions demands reasoning about the complex and continuous dynamics of the cloth. These challenges underscore the need for advancements in (i) state estimation and (ii) dynamics modeling for robust cloth manipulation.
@@
State estimation for cloth is particularly difficult due to its near-infinite degrees of freedom and frequent self-occlusions. While humans intuitively infer full object shapes from partial observations using prior experience, most existing methods are unable to fully capture the complex mapping between highly partial observations and high-dimensional object states [7, 15, 1]. A promising direction is to develop perception models that “imagine” full states from partial observations by leveraging extensive prior experience, akin to human reasoning.
Modeling cloth dynamics presents additional challenges due to the object’s high degrees of freedom and nonlinear behavior. Current approaches often use particle- or mesh-based representations and employ graph neural networks (GNNs) for state transition modeling [44, 15, 18]. GNNs offer advantages in data-scarce domains through their inductive biases, such as spatial equivariance and locality. However, even with locally constructed graphs to reduce computational overhead, the number of edges can grow substantially with the number of nodes, posing scalability issues [30]. Additionally, the locality of graph structures often limits their ability to propagate information effectively over long distances, which is critical for accurately modeling cloth dynamics. Thus, there is a need for more expressive and scalable alternatives.
@@
This work formulates state estimation and dynamics prediction as conditional generation tasks. State estimation reconstructs full states from partial observations, while dynamics prediction generates future states conditioned on the current state and robot actions. To model these complex high-dimensional mappings, we employ diffusion-based models, inspired by their recent successes in computer vision [22, 19], science [31], and robotics [8]. We hypothesize that diffusion models enable accurate state reconstruction and precise trajectory synthesis. Additionally, we adopt Transformer architectures [34] instead of GNNs to overcome locality constraints, leveraging self-attention to capture both local and global dependencies in deformable object dynamics.
@@
Building on these insights, we introduce UniClothDiff, a unified framework comprising a Diffusion Perception Model (DPM), a Diffusion Dynamics Model (DDM), and a model-predictive controller for cloth manipulation. The DPM employs diffusion models and Transformers to generate full-state reconstructions from potentially highly occluded observations, significantly outperforming prior methods. The DDM uses the same architectural principles to achieve long-horizon dynamics prediction, reducing error rates by an order of magnitude compared to GNN-based approaches. Our models are trained on a large-scale dataset with 200K samples generated in the simulator and demonstrate zero-shot transfer to real-world cloth manipulation. Furthermore, DDM is designed with an embodiment-agnostic action representation, enabling deployment on different robotic platforms, including a parallel gripper and a dexterous hand. Extensive experiments show superior system performance compared to earlier methods.
II Related Work
II-A Deformable Object Manipulation
@@
The manipulation of deformable objects, particularly garments, remains a fundamental challenge in robotics due to high-dimensional state spaces and complex dynamics. Current approaches fall into two main categories: model-free and model-based methods. Model-free approaches, including reinforcement learning (RL) [23, 16] and imitation learning (IL) [3, 10, 8, 43, 26], learn direct observation-to-action mappings through end-to-end training. However, these methods struggle with precise shape control due to the lack of explicit object dynamics modeling. Model-based approaches require accurate state estimation [6, 32, 33, 2, 14], challenging when deformable objects like cloth exhibit severe self-occlusions. Furthermore, learning dynamics models demands extensive training data covering large state and action spaces. To overcome these challenges, we propose leveraging the expressive generative models, specifically diffusion models, for both full-state estimation from partial observations and dynamics modeling using large-scale simulation data.
II-B Learning-Based Dynamics Models
@@
Learning-based dynamics models aim to predict state transitions directly from interaction data, with the choice of state representation playing a critical role. Pixel-based representations frame the problem as action-conditioned video prediction [12, 40]. However, these models often require extensive training data, are sensitive to occlusions, and struggle to generate physically plausible predictions due to the lack of explicit 3D structure [41]. In contrast, particle-based representations, like point clouds and meshes, offer a more structured and physically grounded approach to modeling deformable objects. These representations are typically processed using graph neural networks (GNNs), which model state transitions through message passing [44, 15, 32, 33, 2]. While GNNs have demonstrated effectiveness, we find that diffusion models provide a more scalable and expressive alternative, enabling accurate state transition learning from large-scale datasets and advancing deformable object dynamics modeling.
II-C Diffusion Models
Diffusion models [11] have emerged as a powerful paradigm in generative modeling, capable of capturing complex, high-dimensional data distributions precisely. They have demonstrated significant success across various domains, including image generation [29, 25], video generation [35, 22], 3D shape synthesis [27], and robotic policy learning [8, 43]. In this study, we extend the representational capabilities of diffusion models to tackle key challenges in deformable object manipulation: perception and dynamics modeling. Specifically, we employ diffusion models to recover full geometric details from partial observations, enabling accurate state estimation despite severe occlusions and complex object configurations. Additionally, we employ diffusion models to learn expressive dynamics models that capture highly non-linear state transitions, enabling robust model-based planning.
III Method
III-A Overview
@@
Figure 2: Overview. (a) Perception: Our Diffusion Perception Model (DPM) reconstructs the full cloth state from a partial point cloud. Using a denoising process parameterized by $\boldsymbol{\epsilon_{\theta}^{p}}$, DPM refines the cloth state over $K$ denoising steps, starting from random noise. (b) Dynamics Prediction: Our Diffusion Dynamics Model (DDM) generates future cloth states based on the current estimated state and robot actions, using a transformer-based architecture. The expressive power of diffusion generative models allows our method to handle severe occlusions in perception and capture the intricate dynamics of cloth manipulation.
@@
At each timestep, the system processes multiview RGB-D observations $o_{t}\in\mathcal{O}$, represented as $o_{t}=\{I_{t}^{0},I_{t}^{1},\ldots,I_{t}^{l-1}\}$ with $l$ camera views, to estimate the cloth’s 3D state $s_{t}\in\mathcal{S}$ given canonical state of the template mesh $s_{c}$. The state of the cloth is defined by a mesh $s_{t}=\{V_{t},E_{t}\}$, where $E_{t}$ represents the invariant edge connectivity and $V_{t}\in\mathbb{R}^{N_{v}\times 3}$ denotes the positions of vertices in 3D space where $N_{v}$ denotes the number of vertices. We propose that generative models can effectively infer unobserved patterns in raw observations, enabling robust state estimation.
Given the estimated state, a learned dynamics model $f$ predicts the future state $s_{t+1}\in\mathcal{S}$ based on state history $s_{t:t-i}\in\mathcal{S}$ and planned action $a_{t}\in\mathcal{A}$. This dynamics model is incorporated into a model-predictive control framework to optimize action sequences for achieving target state $s_{g}$:
@@ III-B State Estimation
We first address the challenging problem of inferring complete cloth configurations from partial observations. Despite using four multi-view RGB-D cameras, accurate state estimation remains difficult due to inherent self-occlusions and complex deformations. Inspired by humans’ capability of imagining full object states from highly partial observations, we propose to use diffusion models, which are expressive generative models that can capture complex data distributions, to recover full states of clothes from observations.
Conditional Diffusion Process
@@
We formulate cloth state estimation as a conditional denoising diffusion process. We use object point cloud as conditioning inputs to the model, as the sim-to-real gap is minimized in the particle space [5].
@@
Specifically, we model the conditional distribution $p(s|s_{c},\mathbf{e}_{\text{pc}})$ using standard denoising diffusion probabilistic model (DDPM) [11], where $s_{c}$ represents the state of the canonical cloth mesh and $\mathbf{e}_{\text{pc}}$ denotes the embedding of the conditional point cloud. To get point cloud embedding, we partition the point cloud into patches by first sampling M center points using farthest point sampling (FPS) and performing K-Nearest Neighbors (KNN) clustering. Then each resulting patch is processed through a PointNet [28] to obtain its embedding representation $\mathbf{e}_{\text{pc}}\in R^{B\times M\times D_{1}}$. where $B$ is batch size and $D_{1}$ is the dimension of point cloud embedding.
In the forward process, starting from $s_{0}$ from the conditional distribution, gaussian noise is gradually added at levels $t\in\{1,...,T\}$ to get noisy state as:
$s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ (1)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$, $\bar{\alpha}_{t}:=\prod_{s=1}^{t}1-\beta_{s}$, and $\{\beta_{1},\ldots,\beta_{T}\}$ is the variance schedule of a process with $T$ steps. In the reverse process, starting from noisy state $s_{t}$ sampled from the normal distribution, the conditional denoising network $\epsilon^{p}_{\theta}$ gradually denoising from $s_{t}$ to $s_{t-1}$ and finally construct $s_{0}$.
Model Architecture
@@
In terms of model design, we adopt vanilla Vision Transformer (ViT) architecture [9] as our backbone. Recent works in image and video generation [25, 22] has demonstrated the scalability of ViT-based diffusion models. The model accepts a point cloud and a canonical mesh as inputs, in addition to the noisy mesh state that requires denoising. First, we describe our input tokenization approach. Then, we detail how condition information is integrated into the model architecture. Finally, we present the training objective.
@@
Tokenization. We tokenize the input mesh as non-overlapping vertex patches. We first use farthest point sampling (FPS) to sample a fixed number of points as patch centers $C\in\mathbb{R}^{N\times 3}$. To patchify the mesh vertices, We use the $N$ centers obtained from FPS to construct a Voronoi diagram in the 3D points space. This tessellation divides the point cloud into N distinct regions, where each region contains all points closer to its associated center than to any other center. Formally, for a set of centers $C={c_{1},c_{2},\ldots,c_{n}}$, the Voronoi cell $V_{i}$ corresponding to center $C_{i}$ is defined as: where $d(x,y)$ denotes the distance between points x and y. Each Voronoi cell is treated as a distinct patch, encompassing a local neighborhood of points. Each patch will go through a PointNet [28] layer for feature extraction. To add positional encoding to each patch, we compute its center in the canonical space. The 3D coordinates $(x,y,z)$ of this center serve as the positional encoding for the entire patch directly which is similar to most point cloud transformer works [37].
@@
Conditioning. Following the tokenization process, the input token is directly subjected to a sequence of transformer blocks for processing. To effectively condition the point cloud embedding, we adopt two approaches. First, The conventional layer normalization is replaced with an adaptive layer normalization (AdaLN)  [39] to better incorporate conditional information. AdaLN modulates the normalization parameters based on the point cloud condition embedding. Specifically, given the tokenized input $x\in\mathbb{R}^{B\times N\times D_{2}}$ and a condition embedding $\textbf{e}_{pc}\in\mathbb{R}^{B\times M\times D_{1}}$, AdaLN first computes the scale $\gamma$ and shift $\beta$ parameters through a learned affine transformation:$[\gamma,\beta]=\textbf{e}_{pc}W_{AdaLn}$ where $W_{AdaLn}\in\mathbb{R}^{D_{1}\times 2D_{1}}$ is a learnable weight matrix. The normalized output is then computed using the learned scale and shift parameters.
Then, we incorporate conditional information through a cross-attention layer positioned after the multi-head self-attention (MHSA). In this cross-attention operation, the hidden states $x$ serve as both query and key vectors, while the conditional information acts as the value vector. The computation proceeds as follows:
$x$ $=\text{CrossAttention}(W_{Q}^{(c)}x,W_{K}^{(c)}\textbf{e}_{pc},W_{V}^{(c)}\textbf{e}_{pc})$ (2)
where $W^{(c)}$ are learnable parameters and the attention operation is computed following standard implementation in  [34]. This adaptive normalization allows the network to dynamically adjust its normalization parameters based on the condition information, enabling more effective feature modulation for different input conditions. The cross-attention mechanism enables effective integration of conditional features into the representation learning process.
@@
Decoding. Finally, the decoding process transforms the hidden states $x$ into 3D vertex coordinates through a two-stage process. First, we employ distance-weighted interpolation to upsample the hidden states, where interpolation weights are computed from canonical-space distances between vertices and their corresponding patch centers. This operation produces an intermediate representation $x\in\mathbb{R}^{B\times N_{v}\times D_{2}}$. A Multi-Layer Perceptron (MLP) then maps this representation to the final output $x_{out}\in\mathbb{R}^{B\times N_{v}\times 3}$, yielding the predicted noise added onto the 3D coordinates for each vertex during the diffusion forward process as illustrated in Equation 1. Details of our model are presented in Appendix B-A.
Training
@@
For state estimation, the denoising model $\epsilon^{p}_{\theta}(s^{(k)}|s_{c},\textbf{e}_{\text{pc}})$ for DPM is trained by minimizing the loss:
$\mathcal{L}_{\text{MSE}}=\left\|\epsilon-\epsilon^{p}_{\theta}\left(\sqrt{1-\beta^{(k)}}s+\sqrt{\beta^{(k)}}\epsilon\middle|s_{c},\textbf{e}_{\text{pc}}\right)\right\|^{2}$
where $\epsilon\sim\mathcal{N}(0,I)$ and $\beta^{(k)}\in\mathbb{R}$ are $K$ different noise levels for $k\in[1,K]$. Training details are presented in Appendix B-B.
III-C Dynamics Prediction
Given the estimated state, the goal of dynamics prediction is to reason about future states of the cloth given robot actions.
Conditional Diffusion Process
To learn the conditional posterior distribution $p(s_{t+1:t+j+1}|a_{t},s_{t:t-h})$, we parameterize it using diffusion models. Here, $a_{t}$ represents the robot action, which is the delta end-effector position, $s_{t:t-i}$ denotes the historical states, and $s_{t+1:t+j+1}$ is the $j$ frame future states to be predicted at timestep $t$. Following prior work  [44, 41], we heuristically set $i=3$ and $j=5$. The diffusion reverse process construct $s_{t}$ conditioned on history frames and action by gradually denoised from normal distribution with the denoising network $\epsilon_{\theta}^{d}$. Since we use delta end-effector position as action representation, to effectively encode the action space, we employ a Fourier feature-based embedding. This approach is particularly well-suited for representing continuous spatial information, such as the 3D coordinates of action points. Given the input action $a\in\mathbb{R}^{B\times N\times 3}$, where $N$ is the number of actions and $3$ represents the dimension of $(x,y,z)$ coordinates, we compute the embedding $\mathbf{e}\in\mathbb{R}^{B\times N\times D_{3}}$ as:
where $D_{3}$ is the action embedding dimension, $d\in\{0,\ldots,D_{3}/6-1\}$, and $f_{d}=100^{d/(D_{3}/6)}$ are the Fourier feature frequencies. The resulting embedding $e$ provides a rich, high-dimensional representation of the action space, which is later injected as the condition into our model through cross-attention layers.
Model Architecture
We extend our state estimation architecture to model dynamics by modifying the condition input to incorporate robot actions and enhancing the temporal modeling capability. Specifically, we augment the spatial attention-focused ViT backbone with dedicated temporal transformer blocks to capture dynamics effectively. The model receives input $x\in\mathbb{R}^{B\times F\times N_{v}\times C}$, where $F$ consecutive frames are concatenated along the temporal dimension. Following the DPM’s tokenization procedure, the basic transformer block generates hidden states $h_{4}\in\mathbb{R}^{(B\times F)\times N\times C}$, which are then processed by the temporal transformer block. To enable temporal processing, we reshape the hidden states from shape $\mathbb{R}^{(B\times F)\times N\times C}$ to $\mathbb{R}^{(B\times N)\times F\times C}$. We incorporate temporal information through positional embeddings $e_{t}$, computed using sinusoidal encoding:
$\omega_{i}=\frac{1}{10000^{2i/C}},\quad i\in[0,C/2-1]$
$e_{t}=[\sin(\mathbf{p}\otimes\boldsymbol{\omega});\cos(\mathbf{p}\otimes\boldsymbol{\omega})]$
where $\mathbf{p}\in\mathbb{R}^{F}$ represents temporal positions and $\otimes$ denotes the Kronecker product.
Training
The denoising model $\epsilon^{d}_{\theta}(s_{t+1}^{(k)}|s_{t:t-i},a_{t})$ for DDM is trained by minimizing loss:
@@ III-C Dynamics Prediction
$\mathcal{L}_{\text{MSE}}=\left\|\epsilon-\epsilon_{\theta}^{d}\left(\sqrt{1-\beta^{(k)}}s_{t+1}+\sqrt{\beta^{(k)}}\epsilon\middle|s_{t:t-i},a_{t},\right)\right\|^{2}$
@@
where $\epsilon\sim\mathcal{N}(0,I)$ and $\beta^{(k)}\in\mathbb{R}$ are $K$ different noise levels for $k\in[1,K]$
III-D Model-Based Planning

### v2

@@ Abstract
Cloth manipulation is challenging due to its highly complex dynamics, near-infinite degrees of freedom, and frequent self-occlusions, which complicate both state estimation and dynamics modeling. Inspired by recent advances in generative models, we hypothesize that these expressive models can effectively capture intricate cloth configurations and deformation patterns from data. Therefore, we propose a diffusion-based generative approach for both perception and dynamics modeling. Specifically, we formulate state estimation as reconstructing full cloth states from partial observations and dynamics modeling as predicting future states given the current state and robot actions. Leveraging a transformer-based diffusion model, our method achieves accurate state reconstruction and reduces long-horizon dynamics prediction errors by an order of magnitude compared to prior approaches. We integrate our dynamics models with model predictive control and show that our framework enables effective cloth folding on real robotic systems, demonstrating the potential of generative models for deformable object manipulation under partial observability and complex dynamics.
Keywords: Deformable Object Manipulation, Dynamics Model Learning, State Estimation, Generative Models, Cross-Embodiment Generalization
1 Introduction
@@
Textile deformable objects, such as clothing, are ubiquitous in daily life. Yet, manipulating these objects is a long-standing challenge in robotics [1, 2], due to their complex geometric structures and dynamics. Effective cloth manipulation requires accurately estimating the state of cloth despite severe self-occlusions, as well as reasoning over its complex, continuous dynamics to optimize actions. These difficulties highlight the need for advancements in both (i) state estimation and (ii) dynamics modeling to enable robust robotic cloth manipulation.
@@
State estimation for cloth is particularly challenging due to frequent self-occlusions arising from its highly deformable structure. While humans intuitively infer full object shapes from partial observations using prior experience, most existing methods are unable to fully capture the complex mapping between highly partial observations and high-dimensional object states [3, 4, 5]. A promising direction is to develop perception models that can “imagine” full states from partial observations by leveraging extensive prior experience, akin to human reasoning.
Modeling cloth dynamics poses another significant challenge due to its highly nonlinear nature. Current approaches typically represent cloth using particle- or mesh-based structures and model their interactions with graph neural networks (GNNs) [6, 4, 7, 8]. GNNs offer advantages in data-scarce domains through spatial equivariance and locality, but they scale inefficiently with the number of graph nodes [9]. Moreover, the locality inherent to graph structures often limits their ability to capture long-range dependencies, which is crucial for accurate dynamics modeling.
@@
In this work, we formulate state estimation and dynamics prediction as conditional generation processes. State estimation reconstructs full states from partial observations, while dynamics prediction generates future states conditioned on the current state and robot actions. To model these complex high-dimensional mappings, we employ diffusion-based models, inspired by their recent successes in capturing complex data distributions in computer vision [10, 11], science [12], and robotics [13]. We hypothesize that diffusion models with scalable architecture (e.g., Transformer [14]) can enable accurate state reconstruction and dynamics modeling.
@@
Building on these insights, we introduce UniClothDiff, a unified framework that integrates a Diffusion Perception Model (DPM), a Diffusion Dynamics Model (DDM), and model predictive control for cloth manipulation. Conceptually, DPM leverages diffusion models and Transformers to reconstruct full cloth states from sparse and occluded RGB-D observations, while DDM predicts long-horizon dynamics conditioned on current states and actions. Trained on a large-scale cloth interaction dataset with 500K transitions in simulation and evaluated in both simulation and real-world, our models achieve substantial performance gains: DPM achieves superior performance compared to prior approaches in cloth state estimation, and DDM reduces long-horizon prediction error by an order of magnitude compared to GNN-based baselines. With an embodiment-agnostic action representation, our framework can be deployed on both parallel grippers and dexterous hands. Real-world experiments demonstrate superior manipulation performance over previous approaches, highlighting the potential of generative modeling in deformable object manipulation.
2 Related Work
@@
Deformable Object Manipulation. Manipulating deformable objects such as garments remains a long-standing challenge in robotics, due to their high-dimensional state space and complex, nonlinear dynamics. Model-free approaches, including reinforcement learning (RL) [15, 16] and imitation learning (IL) [17, 18, 13, 19, 20], learn direct observation-to-action mappings through end-to-end training. However, these methods struggle with precise shape control due to the lack of explicit dynamics reasoning. Model-based approaches require accurate state estimation [21, 22, 23, 24, 25, 26], which is highly challenging with partial observations. Further, learning dynamics models demands extensive training data to cover large state and action spaces. Thus, we propose to learn expressive generative models for state estimation and dynamics modeling using large-scale simulation data.
@@
Learning-Based Dynamics Models. Learning-based dynamics models [27] predict state transitions from interaction data, where the choice of state representation is crucial. Pixel-based models view the problem as action-conditioned video prediction [28, 29], but they are often sample-inefficient, vulnerable to occlusions, and lack physical realism for contact-rich scenarios [30]. Structured representations, such as particles or meshes, provide stronger physical priors and are typically coupled with graph neural networks (GNNs) that perform inference via message passing [24, 6, 4, 22, 23, 8]. While sample-efficient, GNNs often struggle with scalability and long-range interactions. In contrast, we find diffusion models offer greater expressiveness and scalability, enabling accurate dynamics prediction from large-scale data and improving modeling of deformable object behavior.
Diffusion Models. Diffusion models [31], a class of generative models with expressive capability of capturing complex, high-dimensional data distributions precisely, have emerged as a powerful paradigm and been applied across diverse domains, including generation of images [32, 33], videos [34, 10], 3D shapes [35], as well as robot policy learning [13, 19] and world modeling [36, 37]. In this work, we adapt diffusion models for deformable objects manipulation, leveraging their superior data distribution modeling capability for (i) estimating full cloth configurations from partial point cloud observations, and (ii) modeling state transitions to enable accurate future prediction and model-based planning for cloth manipulation.
3 Method
3.1 Overview
@@
Figure 1: Overview. (a) Perception: Our Diffusion Perception Model (DPM) reconstructs the full cloth state from a partial point cloud. Using a denoising process parameterized by $\boldsymbol{\epsilon_{\theta}^{p}}$, DPM refines the cloth state over $K$ denoising steps, starting from random noise. (b) Dynamics Prediction: Our Diffusion Dynamics Model (DDM) generates future cloth states based on the current estimated state and robot actions, using a transformer-based architecture.
@@
At each timestep, the system processes multiview RGB-D observations $o_{t}\in\mathcal{O}$, represented as $o_{t}=\{I_{t}^{0},I_{t}^{1},\ldots,I_{t}^{l-1}\}$ with $l$ camera views, to estimate the cloth’s 3D state $s_{t}\in\mathcal{S}$ given canonical state of the template mesh $s_{c}$. The state of the cloth is defined by a mesh $s_{t}=\{V_{t},E_{t}\}$, where $E_{t}$ represents the invariant edge connectivity and $V_{t}\in\mathbb{R}^{N_{v}\times 3}$ denotes the positions of vertices in 3D space where $N_{v}$ denotes the number of vertices. We propose that generative models can effectively infer unobserved patterns in partial RGB-D observations, enabling robust state estimation.
Given the estimated state, a learned dynamics model $f$ predicts the future state $s_{t+1}\in\mathcal{S}$ based on state history $s_{t-i:t}\in\mathcal{S}$ and planned action $a_{t}\in\mathcal{A}$. This dynamics model is integrated with model-predictive control to optimize action sequences for achieving the target state $s_{g}$:
@@ 3.2 State Estimation
We first address the problem of inferring cloth configurations from partial observations. Despite using four multi-view RGB-D cameras, severe self-occlusions make accurate state estimation infeasible. Inspired by the human ability to infer hidden object states from partial views, we propose using diffusion models to generate full cloth configurations from limited observations.
@@
Conditional Diffusion Process. We formulate cloth state estimation as a conditional denoising diffusion process, using the object point cloud as the conditioning input. Conditioning on point clouds helps minimize the sim-to-real gap due to their nature as a mid-level visual representation and maintains geometric invariance [38, 39].
@@
Specifically, we model the conditional distribution $p(s|s_{c},\mathbf{e}_{\text{pc}})$ using standard denoising diffusion probabilistic model (DDPM) [31], where $s_{c}$ represents the state of the canonical cloth mesh and $\mathbf{e}_{\text{pc}}$ denotes the embedding of the conditional point cloud. To get point cloud embedding, we partition the point cloud into patches by first sampling M center points using farthest point sampling (FPS) and performing K-Nearest Neighbors (KNN) clustering. Then each resulting patch is processed through a PointNet [40] to obtain its embedding representation $\mathbf{e}_{\text{pc}}\in R^{B\times M\times D_{1}}$. where $B$ is the batch size and $D_{1}$ is the dimension of the point cloud embedding.
[…]
@@
Model Architecture. We adopt vanilla Vision Transformer (ViT) architecture [41] as our backbone, which has been shown to be highly scalable in image and video generation [33, 10]. The model takes a point cloud and a canonical template mesh as input, in addition to the noisy mesh state that requires denoising. We detail our network architecture and training objective below.
@@
Tokenization. We tokenize the input mesh as non-overlapping vertex patches in canonical space. We first use farthest point sampling (FPS) to sample a fixed number of points as patch centers $C\in\mathbb{R}^{N\times 3}$. To patchify the mesh vertices, we use the $N$ centers obtained from FPS to construct a Voronoi diagram in the 3D points space. This tessellation divides the point cloud into N distinct regions, where each region contains all points closer to its associated center than to any other center. Each Voronoi cell is treated as a distinct patch, encompassing a local neighborhood of points which will then go through a PointNet [40] layer for feature extraction.
@@
Conditioning. Following the tokenization process, the input token is directly subjected to a sequence of transformer blocks for processing. To effectively condition the point cloud embedding, we adopt two approaches. First, the conventional layer normalization is replaced with an adaptive layer normalization (AdaLN) [42] to better incorporate conditional information, which modulates the normalization parameters based on the point cloud condition embedding for effective feature modulation. Then, we incorporate conditional information through a cross-attention layer positioned after the multi-head self-attention (MHSA). In this cross-attention operation, the hidden states $x$ serve as the query vector, while the conditional information acts as both the key and value vectors. The computation proceeds as $x=\text{CrossAttention}(W_{Q}^{(c)}x,W_{K}^{(c)}\textbf{e}_{pc},W_{V}^{(c)}\textbf{e}_{pc})$ where $W^{(c)}$ are learnable parameters, enabling effective conditioning during the learning process.
@@
Decoding. Finally, the decoding process transforms the hidden states $x$ into 3D vertex coordinates through a two-stage process. First, we employ distance-weighted interpolation to upsample the hidden states, where interpolation weights are computed from canonical-space distances between vertices and their corresponding patch centers. This operation produces an intermediate representation $x\in\mathbb{R}^{B\times N_{v}\times D_{2}}$. A Multi-Layer Perceptron (MLP) then maps this representation to the final output $x_{out}\in\mathbb{R}^{B\times N_{v}\times 3}$, yielding the predicted noise added onto the 3D coordinates for each vertex during the diffusion forward process. Details of our model are presented in Appendix C.2.
@@
Training. We train the denoising model $\epsilon^{p}_{\theta}(s^{(k)}|s_{c},\textbf{e}_{\text{pc}})$ to minimize the loss:
@@
$\mathcal{L}(\theta)=\mathbb{E}_{s,s_{c},\mathbf{e}_{\text{pc}}\sim p_{\text{data}}}\left[\left\|\epsilon-\epsilon^{p}_{\theta}\left(\sqrt{1-\beta^{(k)}}s+\sqrt{\beta^{(k)}}\epsilon\middle|s_{c},\mathbf{e}_{\text{pc}}\right)\right\|^{2}\right]$
@@
where $\epsilon\sim\mathcal{N}(0,I)$ and $\beta^{(k)}\in\mathbb{R}$ are $K$ different noise levels for $k\in[1,K]$. Details of the training process are available in Appendix C.3.
3.3 Dynamics Prediction
Given the estimated state, the goal of dynamics prediction is to reason about future states of the cloth given robot actions. We extend our state estimation architecture to model dynamics by modifying the condition input to incorporate robot actions and enhancing the temporal modeling capability with additional temporal attention layers. The remaining components, including tokenization, training objective, and decoding of the model, are identical to those in the state estimation framework.
Conditional Diffusion Process. To learn the conditional posterior distribution $p(s_{t+1:t+j+1}|a_{t},s_{t-i:t})$, we parameterize it using diffusion models. Here, $a_{t}$ represents the robot action, $s_{t-i:t}$ denotes the historical states, and $s_{t+1:t+j+1}$ is the $j$ frame future states to be predicted at timestep $t$. Following prior work  [6, 30], we heuristically set $i=3$ and $j=5$. The diffusion reverse process construct $s_{t}$ conditioned on history frames and action by gradually denoising from a normal distribution with the denoising network $\epsilon_{\theta}^{d}$. Since we use delta end-effector position as action representation, to effectively encode the action space, we employ a Fourier feature-based embedding following NeRF [43] to represent continuous spatial information, with detailed formulation in Appendix C.2.
3.4 Model-Based Planning